# JDT3D: Addressing the Gaps in LiDAR-Based Tracking-by-Attention

### v1

@@ Abstract
Multi-Object Tracking Autonomous Vehicles
@@ 1 Introduction
One insight is that LiDAR-based tracking methods usually suffer from sparse supervision signals due to point cloud and object sparsity. Inspired by object sampling augmentation [25] for LiDAR-based detection methods, we propose track sampling, a temporally consistent augmentation method that injects consistent objects over multiple LiDAR frames to enrich supervision signals while maintaining temporal consistency.
Also, existing TBA methods usually maintain an inconsistency between training and inference regarding their query propagation method. During training, queries that are matched to ground truth samples are propagated to the next frame, but at inference, propagation is based on a confidence threshold. This inconsistency between training and inference could confuse the model to over-trust false positive queries. To address this, we propose a confidence threshold propagation strategy consistent for both training and inference and conduct comprehensive evaluations and discussions on the strategy design.
@@ 2.2 Joint Detection and Tracking (JDT)
Joint detection and tracking (JDT) is a recent paradigm in both 2D and 3D MOT that involves training a model to perform detection and tracking end-to-end. In contrast to the TBD methods that rely on a dedicated detector to generate detections, JDT methods simultaneously output object detections with associated tracking IDs. Without heuristic matching or hand-crafted rules, SimTrack [16] builds upon CenterPoint [26] by learning a hybrid-time centerness map between two consecutive LiDAR point clouds. AlphaTrack [28] adds a cross-modal fusion scheme to the CenterPoint tracking method[26] and applies an alternating training strategy to balance between the detection branch and appearance affinity. The above two methods perform detection and tracking jointly, but the tracks are not utilized to inform future detections. 3D DetecTrack [9] proposes tracklet-aware 3D detection that reconfigures the initial detection using the latest tracklets’ information and employs a spatio-temporal gated GNN for association.
Tracking-by-attention   A climbing trend within JDT is the tracking-by-attention (TBA) approach that emerged from TrackFormer [18] for 2D MOT. TBA uses the transformer architecture [22] and formulates tracking as a frame-to-frame set prediction problem that can be trained in an end-to-end manner.
@@ 3 Method
To analyze LiDAR-based TBA, we developed a baseline architecture since no implementations of LiDAR-based TBA methods are available as of the writing of this paper. We build on aspects from query-based 3D object detection [1] and vision-only TBA [19] to formulate our method, called JDT3D. JDT3D is summarized in Fig. 2.
@@ 4.1 Dataset and Metrics
JDT3D is trained and evaluated on the nuScenes dataset [2], a multi-modal autonomous driving dataset containing data from six multi-view cameras, five radars, and a 360-degree LiDAR. The dataset contains a total of 40,157 frames across 1000 scenes, of which 850 scenes are annotated for training (700) and validation (150) at a frame rate of 2 Hz. We train on the seven object classes annotated for the tracking task: bicycle, motorcycle, pedestrian, bus, car, trailer, and truck.
@@ 4.2 Implementation Details
VoxelNet [25, 32] is used as the LiDAR backbone to extract the BEV feature map. For each frame, the top 200 predictions from the heatmap are used to initialize the proposal queries. The training clip length $L$ is set to 3 frames, limited by available compute. The confidence threshold for track queries $\tau_{pass}$ is set to 0.4 for training and inference. The Hungarian matching cost for ground truth assignment consists of the sum of the L1 cost on the predicted box and the focal loss [14] on the predicted class confidence. All detections are generated in the car ego-frame, so the object centers of the track queries are updated to their relative positions in the current training frame based on the car’s ego-motion between time steps. A pretrained VoxelNet backbone from BEVFusion [15] is used to initialize the feature extractor. To train JDT3D, the model is first trained on the detection task using $L=1$ for 20 epochs. It is then further trained on the tracking task using $L=3$ for another 20 epochs. All experiments were trained and evaluated on eight NVIDIA V100 GPUs using random flipping, rotation, translation, and scaling on the training sequence, and track sampling augmentation, unless otherwise specified. The inference speed of JDT3D on a Tesla V100 GPU is at 3.85 Hz.
@@ 4.3 Main Results
Our main results compare JDT3D with baseline TBD methods and relevant TBA methods evaluated on the nuScenes test set [2], summarized in Tab. 1. JDT3D outperforms the current state-of-the-art LiDAR-based TBA methods, with a relative improvement of 12.5% on the AMOTA metric and 15.5% on the AMOTP metric on the test split. It also outperforms the fusion-based MotionTrack baseline with LiDAR-only input with a relative improvement of 4.36% on the AMOTA metric and 3.79% on the AMOTP metric. While JDT3D underperforms compared to baseline TBD methods on most metrics, it achieves the lowest number of ID switches among LiDAR-only methods, indicating a strong track coherence.
@@ 4.4 Ablation Studies
Track Sampling.   By comparing the performance of JDT3D on the nuScenes validation set with and without track sampling augmentation, we verify that sampling extra tracks into a training sequence significantly improves the convergence rate and the final performance as shown in Fig. 4. The comparison summarized in Tab. 2 shows a 97% relative improvement on the AMOTA metric from track sampling augmentation, with a similar trend across all other metrics, including detection metrics such as the mAP. The benefits of implementing track sampling are expected to generalize to other LiDAR-based TBA methods, but there are currently no available implementations to verify this hypothesis.
Figure 4: A comparison of the AMOTA convergence with and without track sampling during training, evaluated every four epochs on the nuScenes validation set [2], showing that track sampling dramatically improves the convergence and performance of JDT3D.
Track Sampling AMOTA $\uparrow$ AMOTP $\downarrow$ IDS $\downarrow$ mAP $\uparrow$
0.314 1.377 388 0.225
0.621 0.790 150 0.399
Training Clip Length.   The available hardware limited the training clip length to a maximum of three frames. To validate this choice of clip length, the performances of JDT3D models trained with two and three-frame clips were compared and summarized in Tab. 3. Based on this analysis, it is evident that JDT3D benefits from the longer training sequence length. We hypothesize that the training process benefits from the longer-term, temporally consistent information, resulting in improved tracking performance.
Table 3: Ablation on the training sequence length, $L$, reported on the nuScenes validation set. Increasing the sequence length from 2 to 3 frames in training JDT3D results in a significant boost in performance.
@@
Query Propagation Method.   In Tab. 4, we compare different query propagation methods; the first method is passing queries that have been matched to a ground truth [19, 18], and the second is passing based on the $\tau_{pass}$ threshold for high confidence predictions as described in Sec. 3.2. Results show that using confidence as a filtering criterion during training enhances the overall tracking metrics AMOTA and AMOTP by a large margin and significantly reduces the number of false positives and false negatives while increasing the number of true positives. While the ID switch errors are higher, this increase is less than $2\%$ relative to the increase in true positives. Thus, we verify that using confidence as the query propagation criterion is more appropriate for the TBA paradigm.
@@
Comparison with TBD Baseline.   We compare the performance between our TBA framework and baseline TBD methods such as the CenterPoint greedy-based tracking method [26] and SimpleTrack [20]. To compare our tracking method to the TBD trackers independent of detection performance, we pass the detections generated by JDT3D to the tracker and evaluate its tracking performance compared to our joint approach. The results in Tab. 6 demonstrate clear advantages of our joint end-to-end approach that offers temporal consistency and data association and reaffirm that the gap in tracking performance lies in the missed detections.
@@ 4.6 Proposal and Track Query Imbalances
Fig. 5 shows further analysis of the JDT3D outputs. During training, the positively matched proposal queries (untracked objects) have much higher confidence in the first frame compared to subsequent frames. As well, the positively matched track queries (previously tracked objects) tend to have much higher confidence than the proposal queries. This indicates that the model has a strong bias to previously tracked objects. Having low average confidence for proposal predictions leads to more missed objects and fewer initialized tracks, explaining the high number of false negatives indicated in Tab. 1.
Figure 5: The average confidence of positively matched proposal and track predictions over a three-frame training sequence. There are no track queries in the first frame.
@@
Table 8: The distribution of positive and negative matches for proposal queries and track queries in the second and third frames of training. There are no track queries in the first frame.
@@ 5.2 Potential Improvements
Firstly, based on the ablation performed to evaluate the training clip size, we observed that three-frame training significantly improved performance over two-frame training. Training on longer clips may lead to further improvements, although it comes with a proportionally larger compute requirement.
@@ 6 Conclusion
This paper presents JDT3D, a novel LiDAR tracking method that makes significant progress in closing the gap between TBD and JDT methods. JDT3D is trained using track sampling augmentation, a generalizable data augmentation method that injects extra long-term tracks into training samples. We show that track sampling augmentation significantly improves training performance and convergence, and we hope that future LiDAR-based trackers can leverage these findings. As well, we show that the query propagation scheme during training also has a significant impact on the performance, showing that using a confidence threshold to filter low confidence predictions improves performance greatly. Based on our experiments, we hypothesize that the complexity of tracking-by-attention combined with the memory constraints imposed by LiDAR feature extractors limiting the size of transformer decoders is the cause of the drop in performance between single and multi-frame variants of JDT3D. The reduced detection performance results in the observed gap between tracking-by-detection and tracking-by-attention. We propose potential improvements to LiDAR TBA in the form of longer training sequences and enabling larger decoder layers.

### v2

@@ Abstract
Multi-Object Tracking Autonomous Vehicles Computer Vision
@@ 1 Introduction
One insight is that LiDAR-based tracking methods can suffer from sparse supervision signals due to point cloud and object sparsity. Inspired by object sampling augmentation [25] for LiDAR-based detection methods, we propose track sampling, a temporally consistent augmentation method that injects consistent objects over multiple LiDAR frames to enrich supervision signals while maintaining temporal consistency.
In addition, existing TBA methods maintain an inconsistency between training and inference regarding their query propagation method. During training, queries that are matched to ground truth samples are propagated to the next frame, but at inference, propagation is based on a confidence threshold. This inconsistency between training and inference could confuse the model to over-trust false positive queries. To address this, we propose a confidence threshold propagation strategy consistent for both training and inference and conduct comprehensive evaluations and discussions on the strategy design.
@@ 2.2 Joint Detection and Tracking (JDT)
Joint detection and tracking (JDT) is a recent paradigm in both 2D and 3D MOT that involves training a model to perform detection and tracking end-to-end. In contrast to the TBD methods that rely on a dedicated detector to generate detections, JDT methods simultaneously output object detections with associated tracking IDs. Without heuristic matching or hand-crafted rules, SimTrack [16] builds upon CenterPoint [26] by learning a hybrid-time centerness map between two consecutive LiDAR point clouds. AlphaTrack [28] adds a cross-modal fusion scheme to the CenterPoint tracking method[26] and applies an alternating training strategy to balance between the detection branch and appearance affinity. The above two methods perform detection and tracking jointly, but known tracks are not utilized to inform future detections. 3D DetecTrack [9] proposes tracklet-aware 3D detection that reconfigures the initial detection using the latest tracklets’ information and employs a spatio-temporal gated GNN for association.
2.2.1 Tracking-by-Attention (TBA)
A growing trend within JDT is the tracking-by-attention (TBA) approach that emerged from TrackFormer [18] for 2D MOT. TBA uses the transformer architecture [22] and formulates tracking as a frame-to-frame set prediction problem that can be trained in an end-to-end manner.
@@ 3 Method
To perform our analysis, we developed our own baseline LiDAR-based TBA method since no such implementations were made available as of the writing of this paper. We build on aspects from query-based 3D object detection [1] and vision-only TBA [19] to formulate our method, called JDT3D. JDT3D is summarized in Fig. 2.
@@ 4.1 Dataset and Metrics
JDT3D is trained and evaluated on the nuScenes dataset [2], a multi-modal autonomous driving dataset containing data from six multi-view cameras, five radars, and a 360-degree LiDAR. The dataset contains a total of 40,157 frames across 1000 scenes, of which 850 scenes are annotated for training (700) and validation (150) at a frame rate of 2 Hz. We train on the seven object classes annotated for the tracking task.
@@ 4.2 Implementation Details
VoxelNet [25, 32] is used as the LiDAR backbone to extract the BEV feature map. For each frame, the top 200 predictions from the heatmap are used to initialize the proposal queries. The training clip length $L$ is set to 3 frames. The confidence threshold for track queries $\tau_{pass}$ is set to 0.4 for training and inference. The Hungarian matching cost for ground truth assignment consists of the sum of the L1 cost on the predicted box and the focal loss [14] on the predicted class confidence. All detections are generated in the car ego-frame, so the object centers of the track queries are updated to their relative positions in the current training frame based on the car’s ego-motion between time steps. A pretrained VoxelNet backbone from BEVFusion [15] is used to initialize the feature extractor. To train JDT3D, the model is first trained on the detection task using $L=1$ for 20 epochs. It is then further trained on the tracking task using $L=3$ for another 20 epochs. All experiments were trained and evaluated on eight NVIDIA V100 GPUs using random flipping, rotation, translation, and scaling on the training sequence, and track sampling augmentation, unless otherwise specified. JDT3D performs inference at 3.85 Hz on a Tesla V100.
@@ 4.3 Main Results
Our main results compare JDT3D with baseline TBD methods and relevant TBA methods evaluated on the nuScenes test set [2], summarized in Tab. 1. JDT3D outperforms the current state-of-the-art LiDAR-based TBA methods, with a relative improvement of 12.5% on the AMOTA metric and 15.5% on the AMOTP metric on the test split. Moreover, with LiDAR input only, JDT3D also outperforms the fusion-based MotionTrack-LC [29] with a relative improvement of 4.36% on AMOTA and 3.79% on AMOTP on the test set. While JDT3D underperforms compared to baseline TBD methods on most metrics, it achieves the lowest number of ID switches among LiDAR-only methods, indicating a strong track coherence.
Track Sampling Rate AMOTA $\uparrow$ AMOTP $\downarrow$ IDS $\downarrow$ mAP $\uparrow$
0.0 0.314 1.377 388 0.225
0.25 0.600 0.864 210 0.3867
0.50 0.622 0.816 203 0.400
1.0 0.621 0.790 150 0.399
@@ 4.4 Ablation Studies
4.4.1 Track Sampling.
By comparing the performance of JDT3D on the nuScenes validation set with and without track sampling augmentation, we verify that sampling extra tracks into a training sequence significantly improves the convergence rate and the final performance. The comparison summarized in Tab. 2 shows a 97% relative improvement on the AMOTA metric from track sampling augmentation compared to no track sampling, with a similar trend across all other metrics, including detection metrics such as the mAP. We also analyzed the effect of varying the number of objects that are sampled by reducing the sampling frequency to 0.5 and 0.25 of the nominal rate. We observe that the AMOTA and mAP metrics plateau after a sampling rate of 0.5, although the AMOTP and IDS metrics see continued improvement with increased track sampling. The benefits of implementing track sampling are expected to generalize to other LiDAR-based TBA methods, but there are currently no available implementations to verify this hypothesis.
4.4.2 Training Clip Length.
To validate our choice of clip length, the performances of JDT3D models trained with two to four frame sequences were compared and summarized in Tab. 3. The best performance is achieved when trained over $L=3$ frames, with a significant jump when moving from $L=2$ to $L=3$. We hypothesize that the training process benefits from the longer-term, temporally consistent information, resulting in improved tracking performance. Moving from $L=3$ to $L=4$, the performance drops slightly while also significantly increasing the train time and required memory.
Table 3: Ablation on the training sequence length, $L$, reported on the nuScenes validation set.
@@
4 0.599 0.783 173 0.392
4.4.3 Query Propagation Method.
In Tab. 4, we compare different query propagation methods; the first method is passing queries that have been matched to a ground truth [19, 18], and the second is passing based on the $\tau_{pass}$ threshold for high confidence predictions as described in Sec. 3.2. Results show that using confidence as a filtering criterion during training enhances the overall tracking metrics AMOTA and AMOTP by a large margin and significantly reduces the number of false positives and false negatives while increasing the number of true positives. While the ID switch errors are higher, this increase is less than $2\%$ relative to the increase in true positives. Thus, we verify that using confidence as the query propagation criterion is more appropriate for the TBA paradigm.
@@
4.4.4 Comparison with TBD Baseline.
We compare the performance between our TBA framework and baseline TBD methods such as the CenterPoint greedy-based tracking method [26] and SimpleTrack [20]. To compare our tracking method to the TBD trackers independent of detection performance, we pass the detections generated by JDT3D to the tracker and evaluate its tracking performance compared to our joint approach. The results in Tab. 6 demonstrate clear advantages of our joint end-to-end approach that offers temporal consistency and data association and reaffirm that the gap in tracking performance lies in the missed detections.
@@ 4.5 Detection Performance
Table 8: The average confidence of positively matched proposal and track predictions over a three-frame training sequence. There are no track queries in the first frame.
Avg. Confidence Frame 1 Frame 2 Frame 3
Positively Matched Proposals 0.669 0.409 0.387
Track Predictions n/a 0.759 0.760
@@ 4.6 Proposal and Track Query Imbalances
Tab. 8 shows further analysis of the JDT3D outputs. During training, the positively matched proposal queries, or untracked newborn objects, are more confident in the first frame than in subsequent frames. Also, the positively matched track queries, or previously tracked objects, tend to be more confident than the newborn objects. This indicates that the model has a strong bias toward previously tracked objects. This bias could lead to more missed objects and fewer initialized tracks, explaining the high number of false negatives indicated in Tab. 1.
@@
Table 9: The distribution of positive and negative matches for proposal queries and track queries in the second and third frames of training.
@@ 5.2 Potential Improvements
Firstly, based on the ablation performed to evaluate the training clip size, we observed that three-frame training significantly improved performance over two-frame training, although extending to four frames reduced performance slightly. Further exploration of the effect of clip length on tracking performance may yield insights on how to better leverage longer term temporal information.
@@ 6 Conclusion
This paper presents JDT3D, a novel LiDAR tracking method that makes significant progress in closing the gap between TBD and TBA methods. JDT3D is trained using track sampling augmentation, a generalizable data augmentation method that injects extra long-term tracks into training samples. We show that track sampling augmentation significantly improves training performance and convergence, and we hope that future LiDAR-based trackers can leverage these findings. As well, we show that the query propagation scheme during training also has a significant impact on the performance, showing that using a confidence threshold to filter low confidence predictions improves tracking performance greatly. Based on our experiments, we hypothesize that the complexity of tracking-by-attention combined with the memory constraints imposed by LiDAR feature extractors limiting the size of transformer decoders is the cause of the drop in detection performance between single and multi-frame variants of JDT3D. We propose further areas of exploration including exploring different training sequence lengths and enabling more decoder layers.